OpenStudy (anonymous):

How much water must be added to 12 L of a 40% solution of alcohol to obtain a 30% solution? A. .5 L B. 5 L C. .4 L D. 4 L

OpenStudy (mertsj):

x= amount of water to add x+12 = amount of 30% solution 0x= amount of alcohol in x L of water .4(12)= amount of alcohol in 12 L of 40% solution .3(x+12)= amount of alcohol in x+12 L of 30% solution 0x+.4(12)=.3(x+12) 4.8=.3x+3.6 1.2=.3x x=4